A car moves 30km/hr for 10 minutes and 50km/hr for 20 minutes. What is its average speed?
1
Expert's answer
2018-10-05T11:22:48-0400
The average speed is
v_av=L/T.
v_av=(l_1+l_2)/(t_1+t_2 )=(v_1 t_1+v_2 t_2)/(t_1+t_2 ).
v_av=(30 10/60+50 20/60)/(10/60+20/60)=43.3 km/h.